OpenStudy (anonymous):

Subtract. Simplify, if possible 6-y/y-7 minus 8y-7/7-y =

OpenStudy (anonymous):

the solution is \[\frac{48}{y-7}+7\]

OpenStudy (anonymous):

(1-7y)/(7-y)

OpenStudy (anonymous):

I came up with the first answer....why is that??

OpenStudy (anonymous):

evaluate the following \[\frac{(7-y)(6-y)-(y-7)(8y-7)}{(y-7)(7-y)}\]

OpenStudy (anonymous):

that gives \[\frac{-(y-7)(7y-1)}{(y-7)(7-y)}\]

OpenStudy (anonymous):

last step is to factorise & cancel y-7 from top and bottom to give \[\frac{−(7y−1)}{−(y−7)}\]

OpenStudy (anonymous):

And that gives 48/y-7+7 right

OpenStudy (anonymous):

yep :D